OpenStudy (anonymous):

HOW WOULD YOU DO A PROMLEM LIKE 2Y+X=8 Y-2X=1 WHAT Y= WHAT DOES X=

ganeshie8 (ganeshie8):

basically you have two equations : 2y + x = 8 y - 2x = 1

ganeshie8 (ganeshie8):

you could try solving them by substitution method heard of it before ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

so which method are you familiar with ?

OpenStudy (anonymous):

like the elimanation prosess

ganeshie8 (ganeshie8):

i like elimination method too, whats your first step in elimination ?

ganeshie8 (ganeshie8):

2y + x = 8 y - 2x = 1 How about multiplying first equation by 2 ?

ganeshie8 (ganeshie8):

4y + 2x = 16 y - 2x = 1 ------------------

ganeshie8 (ganeshie8):

add them vertically, what do you get ?

OpenStudy (anonymous):

here is a ex:First you'll need to set up your equations from the information in the problem description. You'll need two variables, one to represent the number of pens and the other to represent the number of pencils. You could use P for pens and W for (wooden) pencils. From the description, you can now write two equations. 2W + 3P = $0.78 and 3W + 2P = $0.72 Now you need to solve this system of equations. One way is to use the "elimination" method. In this method, you would multiply one or both equations by the appropriate integers so as to make both equations have the same number of one of the two variables. This sounds more complicated than it really is. Take a look below: 2(2W + 3P = $0/78) = 4W + 6P = $1.56 3(3W + 2P = $0.72) = 9W + 6P = $2.16 Now, you subtract the first new equation from the second new equation and when you do this, you eliminate the variable P, leaving you with just one variable (W), the number of pencils. After the subtraction, you get: 5W = $0.60 Now divide both sides of the equation by 5 to get W. W = $0.12 So, one pencil costs $0.12 or 12 cents.

ganeshie8 (ganeshie8):

Adding them vertically you get : 4y + 2x = 16 y - 2x = 1 ------------------ 5y + 0 = 17 5y = 17 y = ?

TheSmartOne (thesmartone):

What's up with all caps question??

OpenStudy (anonymous):

@cyncyndgjuyt You see, ganeshie did the same thing as your example. Do you understand what is going on here? You need to work with us here so we know where you stand and where exactly you're having problems.